OpenStudy (anonymous):

how to rewrite y=9x+2 for x

OpenStudy (anonymous):

first you subtract 2 from both sides to get y-2=9x then you divide both sides by 9 to get \[(y-2)/9=x\]

OpenStudy (shadowlegendx):

y = 9x + 2 So 9x + 2 = 9x + 2

OpenStudy (shadowlegendx):

Solve for x

OpenStudy (shadowlegendx):

Remember the first step for solving an algebraic equation is to isolate the variable

OpenStudy (anonymous):

shadowlegendx i didnt understand nothing u just typed

OpenStudy (shadowlegendx):

@Mrs.Mckinney you are trying to find x, correct?

OpenStudy (anonymous):

yes i am

OpenStudy (shadowlegendx):

Ok so since y = 9x + 2 9x + 2 = 9x + 2 is a true statement

OpenStudy (shadowlegendx):

*equation

OpenStudy (anonymous):

That doesn't work, 9x+2=9x+2 is equivalent to 0=0, when you are trying to solve for x, that means you are trying to get x on one side

OpenStudy (shadowlegendx):

@kas9987 what would you suggest then? I am running on no sleep so I'm not at my best, haha

OpenStudy (anonymous):

ShadowLegendX so you are saying thats the answer but its not cause your trying to get x by itself

OpenStudy (shadowlegendx):

\[\frac{ y - 2 }{ 9 } = x \] ?

OpenStudy (anonymous):

@ShadowLegendX I posted my answer, it's the very first one on this page

OpenStudy (anonymous):

so kas9987 can u explain to me how to rewrite y=9x+2 for x please

OpenStudy (shadowlegendx):

@Mrs.Mckinney Read his answer at the top

OpenStudy (anonymous):

first you subtract 2 from both sides to get y-2=9x then you divide both sides by 9 to get (y−2)/9=x

OpenStudy (anonymous):

so kas9987 thats the correct answer??

OpenStudy (shadowlegendx):

\[\frac{ y - 2 }{ 9 } = x\]

OpenStudy (shadowlegendx):

Is that one of your answer choices @Mrs.Mckinney ???

OpenStudy (anonymous):

no when u divide 9 from both sides what is the next step

OpenStudy (shadowlegendx):

That's it, you get the answer that I posted above

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

kas9987 thats the correct answer

OpenStudy (anonymous):

@Mrs.Mckinney yes, when you divide by 9, x is by itself on one side, so you are done, and I'm like 99% sure thats the right answer. It could also be written as \[\frac{ y }{ 9 }-\frac{ 2 }{ 9 }= x\]

OpenStudy (anonymous):

can you rewrite the answer i dont undersatnd what u put PLEASE!!!!

OpenStudy (anonymous):

whats the answer to df=g-10 for d

OpenStudy (anonymous):

oh well then just use what i first wrote, and for df=g-10 you just divide by f so you have d= a fraction of g-10 on the top and f on the bottom, or\[d=\frac{ g-10 }{ f }\]